# How Far Are We on the Decision-Making of LLMs? Evaluating LLMs’ Gaming Ability in Multi-Agent Environments

### v2

@@ Abstract
Decision-making, a complicated task requiring various types of abilities, presents an excellent framework for assessing Large Language Models (LLMs). Our research investigates LLMs’ decision-making capabilities through the lens of a well-established field, Game Theory. We focus specifically on games that support the participation of more than two agents simultaneously. Subsequently, we introduce our framework, $\gamma$-Bench, including eight classical multi-agent games. We design a scoring scheme to assess a model’s performance in these games quantitatively. Through $\gamma$-Bench, we investigate LLMs’ robustness, generalizability, and enhancement strategies. Results reveal that while GPT-3.5 shows satisfying robustness, its generalizability is relatively limited. However, its performance can be improved through approaches such as Chain-of-Thought. Additionally, we conduct evaluations across various LLMs and find that GPT-4 outperforms other models on $\gamma$-Bench, achieving a score of 60.5. Moreover, Gemini-1.0-Pro and GPT-3.5 (0613, 1106, 0125) demonstrate similar intelligence on $\gamma$-Bench. The code and experimental results are made publicly available via https://github.com/CUHK-ARISE/GAMABench.
Figure 1: $\gamma$-Bench enables various LLMs and humans to participate in multi-agent, multi-round games. The framework includes eight classical games in Game Theory, each categorized into one of three groups.
@@ 1 Introduction
Table 1: Performance (scores) of different LLMs on $\gamma$-Bench.
$\gamma$-Bench Leaderboard GPT-3.5 GPT-4 Gemini-Pro
0613 1106 0125 0125 1.0
Guess 2/3 of the Average $41.4_{\pm 0.5}$ $68.5_{\pm 0.5}$ $63.4_{\pm 3.4}$ $91.6_{\pm 0.6}$ $77.3_{\pm 6.2}$
El Farol Bar $74.8_{\pm 4.5}$ $64.3_{\pm 3.1}$ $68.7_{\pm 2.7}$ $23.0_{\pm 8.1}$ $33.5_{\pm 10.3}$
Divide the Dollar $42.4_{\pm 7.7}$ $70.3_{\pm 3.3}$ $68.6_{\pm 2.4}$ $98.1_{\pm 1.9}$ $77.6_{\pm 3.6}$
Public Goods Game $38.8_{\pm 8.1}$ $43.5_{\pm 12.6}$ $17.8_{\pm 1.7}$ $89.2_{\pm 1.8}$ $68.5_{\pm 7.6}$
Diner’s Dilemma $67.0_{\pm 4.9}$ $1.4_{\pm 1.3}$ $2.8_{\pm 2.8}$ $0.9_{\pm 0.7}$ $3.1_{\pm 1.5}$
Sealed-Bid Auction $5.5_{\pm 1.0}$ $3.4_{\pm 1.0}$ $4.2_{\pm 0.3}$ $8.7_{\pm 0.7}$ $14.9_{\pm 6.6}$
Battle Royale $19.5_{\pm 7.7}$ $35.7_{\pm 6.9}$ $28.6_{\pm 11.0}$ $86.8_{\pm 9.7}$ $16.5_{\pm 6.9}$
Pirate Game $68.4_{\pm 20.0}$ $69.6_{\pm 14.7}$ $71.6_{\pm 7.6}$ $85.4_{\pm 8.6}$ $57.4_{\pm 14.3}$
Overall $43.5_{\pm 2.2}$ $44.8_{\pm 3.6}$ $41.9_{\pm 2.3}$ $60.5_{\pm 2.7}$ $43.3_{\pm 3.2}$
@@
To develop a framework to evaluate the decision-making ability of LLMs, we draw upon the principles of Game Theory, a well-established field with extensive applicability. Our rationale is threefold: (1) Scope: Game theory allows for the abstraction of diverse real-life scenarios into simple mathematical models, facilitating a broad range of evaluations. (2) Quantifiability: By examining the Nash equilibrium within these models, we gain a measurable metric for comparing LLMs’ decision-making performance. (3) Variability: The adjustable parameters of these models enable the creation of variant scenarios, enhancing the diversity and robustness of our assessments. Our framework evaluates LLMs in complex settings involving multi-player, multi-action, and multi-round games, incorporating eight classical games extensively analyzed in game theory literature.
@@
The first category in our framework evaluates LLMs’ ability to make optimal decisions by understanding game rules and recognizing patterns in other players’ behavior. A distinctive characteristic of these games is that individual players cannot achieve higher gains without cooperation, provided that other participants cooperate. Essentially, these games’ Nash equilibrium aligns with maximizing overall social welfare. We name such games as Cooperative Games, including (1) Guess 2/3 of the Average, (2) El Farol Bar, and (3) Divide the Dollar. The second category assesses the propensity of LLMs to prioritize self-interest, potentially betraying others for greater gains. In contrast to the first category, games in this category incentivize higher rewards for participants who betray their cooperative counterparts. Typically, the Nash equilibrium in these games leads to reduced social welfare. This category is termed Betraying Games, including (4) Public Goods Game, (5) Diner’s Dilemma, (6) Sealed-Bid Auction. Last but not least, we focus specifically on two games characterized by sequential decision-making processes, distinguishing them from the previous six games based on simultaneous decision-making. These Sequential Games are the (7) Battle Royale and (8) Pirate Game.
@@
In this paper, we first instruct ten agents, based on the gpt-3.5-turbo-0125 model, to engage in the eight games in $\gamma$-Bench, followed by an analysis of the results obtained. Subsequently, we assess the model’s robustness against multiple runs, temperature parameter alterations, and prompt template variations. Further exploration is conducted to ascertain if instructional prompts, such as Chain-of-Thought, enhance the model’s decision-making capabilities. Additionally, the model’s capacity to generalize across diverse game settings is examined. Finally, the performance of LLMs, including GPT-3.5 (0613, 1106, 0125), GPT-4 (0125) and Gemini Pro (1.0), is evaluated.
@@
  • This paper reviews and compares existing literature on the evaluation of LLMs using game theory models, emphasizing variations in focus on LLMs, games, and other settings.
@@
  • The $\gamma$-Bench framework is applied to various LLMs, facilitating an in-depth analysis of their performance in multi-agent gaming scenarios.
2 Background
2.1 Game Theory
Formulation

### v3

@@ Abstract
Decision-making, a complicated task requiring various types of abilities, presents an excellent framework for assessing Large Language Models (LLMs). Our research investigates decision-making capabilities of LLMs through the lens of Game Theory. We focus specifically on games that support the simultaneous participation of more than two agents. We introduce GAMA($\gamma$)-Bench, which evaluates LLMs’ Gaming Ability in Multi-Agent environments. $\gamma$-Bench includes eight classical multi-agent games and a scoring scheme specially designed to quantitatively assess LLMs’ performance. Leveraging $\gamma$-Bench, we investigate LLMs’ robustness, generalizability, and strategies for enhancement. Results reveal that while GPT-3.5 shows satisfying robustness, its generalizability is relatively limited. However, its performance can be improved through approaches such as Chain-of-Thought. Additionally, we evaluate twelve versions from six models, including GPT-3.5, GPT-4, Gemini, LLaMA-3.1, Mixtral, and Qwen-2. We find that Gemini-1.5-Pro outperforms other models with a score of $63.8$ out of $100$, followed by LLaMA-3.1-70B and GPT-4 with scores of $60.9$ and $60.5$, respectively. The code and experimental results are made publicly available via https://github.com/CUHK-ARISE/GAMABench.
Figure 1: $\gamma$-Bench enables multiple LLMs and humans to engage in multi-round games. The framework comprises three categories of games, each targeting different LLM abilities, and includes eight classic games from Game Theory.
@@ 1 Introduction
To develop a framework to evaluate the decision-making ability of LLMs, we draw upon the principles of Game Theory, a well-established field with extensive applicability. Our rationale is threefold: (1) Scope: Game theory allows for the abstraction of diverse real-life scenarios into simple mathematical models, facilitating a broad range of evaluations. (2) Quantifiability: By examining the Nash equilibrium Nash (1950) within these models, we gain a measurable metric for comparing LLMs’ decision-making performance. (3) Variability: The adjustable parameters of these models enable the creation of variant scenarios, enhancing the diversity and robustness of our assessments. Our framework evaluates LLMs in complex settings involving multi-player, multi-action, and multi-round games, incorporating eight classical games extensively analyzed in game theory literature.
Table 1: Performance (scores) of different closed-source LLMs on $\gamma$-Bench.
$\gamma$-Bench Leaderboard GPT-3.5 GPT-4 Gemini-Pro
0613 1106 0125 0125 1.0 1.5
Guess 2/3 of the Average $41.4_{\pm 0.5}$ $68.5_{\pm 0.5}$ $63.4_{\pm 3.4}$ $91.6_{\pm 0.6}$ $77.3_{\pm 6.2}$ $95.4_{\pm 0.5}$
El Farol Bar $74.8_{\pm 4.5}$ $64.3_{\pm 3.1}$ $68.7_{\pm 2.7}$ $23.0_{\pm 8.1}$ $33.5_{\pm 10.3}$ $37.2_{\pm 4.2}$
Divide the Dollar $42.4_{\pm 7.7}$ $70.3_{\pm 3.3}$ $68.6_{\pm 2.4}$ $98.1_{\pm 1.9}$ $77.6_{\pm 3.6}$ $93.8_{\pm 0.3}$
Public Goods Game $38.8_{\pm 8.1}$ $43.5_{\pm 12.6}$ $17.8_{\pm 1.7}$ $89.2_{\pm 1.8}$ $68.5_{\pm 7.6}$ $100.0_{\pm 0.0}$
Diner’s Dilemma $67.0_{\pm 4.9}$ $1.4_{\pm 1.3}$ $2.8_{\pm 2.8}$ $0.9_{\pm 0.7}$ $3.1_{\pm 1.5}$ $35.9_{\pm 5.3}$
Sealed-Bid Auction $5.5_{\pm 1.0}$ $3.4_{\pm 1.0}$ $4.2_{\pm 0.3}$ $8.7_{\pm 0.7}$ $14.9_{\pm 6.6}$ $13.3_{\pm 4.6}$
Battle Royale $19.5_{\pm 7.7}$ $35.7_{\pm 6.9}$ $28.6_{\pm 11.0}$ $86.8_{\pm 9.7}$ $16.5_{\pm 6.9}$ $81.3_{\pm 7.7}$
Pirate Game $68.4_{\pm 20.0}$ $69.6_{\pm 14.7}$ $71.6_{\pm 7.6}$ $85.4_{\pm 8.6}$ $57.4_{\pm 14.3}$ $53.7_{\pm 1.2}$
Overall $43.5_{\pm 2.2}$ $44.8_{\pm 3.6}$ $41.9_{\pm 2.3}$ $60.5_{\pm 2.7}$ $43.3_{\pm 3.2}$ $63.8_{\pm 1.6}$
Table 2: Performance (scores) of different open-source LLMs on $\gamma$-Bench.
$\gamma$-Bench Leaderboard LLaMA-3.1 Mixtral Qwen-2
8B 70B 405B 8x7B 8x22B 72B
Guess 2/3 of the Average $85.5_{\pm 3.0}$ $84.0_{\pm 1.7}$ $94.3_{\pm 0.6}$ $91.8_{\pm 0.4}$ $83.6_{\pm 4.6}$ $93.2_{\pm 1.3}$
El Farol Bar $75.7_{\pm 2.2}$ $59.7_{\pm 3.5}$ $20.5_{\pm 24.2}$ $66.8_{\pm 5.8}$ $39.3_{\pm 12.2}$ $17.0_{\pm 25.5}$
Divide the Dollar $56.4_{\pm 8.4}$ $87.0_{\pm 4.1}$ $94.9_{\pm 1.0}$ $1.2_{\pm 2.8}$ $79.0_{\pm 9.6}$ $91.9_{\pm 2.4}$
Public Goods Game $19.6_{\pm 1.0}$ $90.6_{\pm 3.6}$ $96.9_{\pm 0.8}$ $27.7_{\pm 11.7}$ $83.7_{\pm 3.5}$ $81.3_{\pm 5.9}$
Diner’s Dilemma $59.3_{\pm 2.4}$ $48.1_{\pm 5.7}$ $14.4_{\pm 4.5}$ $76.4_{\pm 7.1}$ $79.9_{\pm 5.8}$ $0.0_{\pm 0.0}$
Sealed-Bid Auction $16.9_{\pm 1.8}$ $4.5_{\pm 0.7}$ $4.2_{\pm 1.2}$ $0.8_{\pm 0.4}$ $5.2_{\pm 1.8}$ $0.9_{\pm 0.2}$
Battle Royale $35.9_{\pm 12.1}$ $77.7_{\pm 26.0}$ $92.7_{\pm 10.1}$ $12.6_{\pm 9.5}$ $36.0_{\pm 21.0}$ $81.7_{\pm 9.6}$
Pirate Game $42.3_{\pm 7.8}$ $35.7_{\pm 7.3}$ $38.2._{\pm 10.8}$ $31.8_{\pm 5.0}$ $53.0_{\pm 2.6}$ $53.0_{\pm 1.1}$
Overall $48.9_{\pm 2.9}$ $60.9_{\pm 2.4}$ $57.0_{\pm 4.2}$ $38.6_{\pm 2.3}$ $57.5_{\pm 2.1}$ $52.4_{\pm 3.4}$
@@
We include eights games and divide them into three categories based on their characteristics. The first category in our framework evaluates LLMs’ ability to make optimal decisions by understanding game rules and recognizing patterns in other players’ behavior. A distinctive characteristic of these games is that individual players cannot achieve higher gains without cooperation, provided that other participants cooperate. Essentially, these games’ Nash equilibrium aligns with maximizing overall social welfare. We name such games as Cooperative Games, including (1) Guess 2/3 of the Average, (2) El Farol Bar, and (3) Divide the Dollar. The second category assesses the propensity of LLMs to prioritize self-interest, potentially betraying others for greater gains. In contrast to the first category, games in this category incentivize higher rewards for participants who betray their cooperative counterparts. Typically, the Nash equilibrium in these games leads to reduced social welfare. This category is termed Betraying Games, including (4) Public Goods Game, (5) Diner’s Dilemma, (6) Sealed-Bid Auction. Last but not least, we focus specifically on two games characterized by sequential decision-making processes, distinguishing them from the previous six games based on simultaneous decision-making. Sequential Games are the (7) Battle Royale and (8) Pirate Game.
@@
In this paper, we instruct ten agents, based on the GPT-3.5 (0125) model, to engage in the eight games, followed by an analysis of the results obtained. Subsequently, we assess the model’s robustness against multiple runs, temperature parameter alterations, and prompt template variations. Further exploration is conducted to ascertain if instructional prompts, such as Chain-of-Thought (CoT) Kojima et al. (2022), enhance the model’s decision-making capabilities. Additionally, the model’s capacity to generalize across diverse game settings is examined. Finally, we evaluate the performance of twelve LLMs, including GPT-3.5-Turbo (0613, 1106, 0125) OpenAI (2022), GPT-4-Turbo (0125) OpenAI (2023), Gemini-1.0-Pro Pichai & Hassabis (2023), Gemini-1.5-Pro Pichai & Hassabis (2024), LLaMA-3.1 (8B, 70B, 405B) Dubey et al. (2024), Mixtral (8x7B, 8x22B) Jiang et al. (2024), and Qwen-2-72B Yang et al. (2024).
@@
  • This paper reviews and compares existing literature on the evaluation of LLMs using game theory models, emphasizing focus on multi-player setting and LLMs’ generalizability.
@@
  • The $\gamma$-Bench framework is applied to twelve LLMs, facilitating an in-depth analysis of their performance in multi-agent gaming scenarios.
2 Games in GAMA-Bench
We collect eight games well studied in Game Theory and propose $\gamma$-Bench, a framework with multi-player, multi-round, and multi-action settings. Notably, $\gamma$-Bench allows the simultaneous participation of both LLMs and humans, enabling us to evaluate LLMs’ performance when playing against humans or fixed strategies. The remaining part of this section details each game.
2.1 Cooperative Games
(1) Guess 2/3 of the Average
Initially introduced by Ledoux (1981), the game involves players independently selecting an integer between 0 and 100 (inclusive). The winner is the player(s) choosing the number closest to two-thirds of the group’s average. A typical initial strategy might lead players to assume an average of 50, suggesting a winning number around $50\times\frac{2}{3}\approx 33$. However, if all participants adopt this reasoning, the average shifts to 33, thereby altering the winning number to approximately 22. The game has a PSNE where all players selecting zero results in a collective win.
(2) El Farol Bar
(3) Divide the Dollar
2.2 Betraying Games
(4) Public Goods Game
(5) Diner’s Dilemma
This game is the multi-player variant of the Prisoner’s Dilemma Glance & Huberman (1994). The game involves $N$ players dining together, with their decision to split all the costs. Each player needs to independently choose whether to order the expensive or the cheap dish, priced at $x$ and $y$ ($x>y$), respectively. The expensive offers $a$ utility per individual, surpassing the $b$ utility of another choice ($a>b$). The game satisfies two assumptions: (1) $a-x<b-y$: Although the expensive dish provides a greater utility, the benefit does not justify its higher cost, leading to a preference for the cheap one when dining alone. (2) $a-\frac{x}{N}>b-\frac{y}{N}$: Individuals are inclined to choose the expensive dish when the cost is shared among all diners. The assumptions lead to an NE where all players opt for the more expensive meal. However, this PSNE results in a lower total social welfare of $N(a-x)$ compared to $N(b-y)$, which is the utility if all choose the cheap one. This game evaluates the long-term perspective and the capacity to establish sustained cooperation.
(6) Sealed-Bid Auction
2.3 Sequential Games
(7) Battle Royale
(8) Pirate Game
This game is a multi-player version of the Ultimatum Game Goodin (1998); Stewart (1999). Each player is assigned a “pirate rank”, determining their action order. The game involves $N$ pirates discussing the division of $G$ golds they have discovered. The most senior pirate first proposes a distribution method. If the proposal is approved by at least half of the pirates, including the proposer, the game ends, and the gold is distributed as proposed. Otherwise, the most senior pirate is thrown overboard, and the next in rank assumes the proposer role until the game ends. Each pirate’s objectives are prioritized as (1) survival, (2) maximizing their share of gold, and (3) the opportunity to eliminate others from the game. Stewart (1999) identifies the optimal strategy, where the most senior pirate allocates one gold to each odd-ranked pirate and keeps the remainder.
3 Vanilla Experiments
This section presents experiments conducted using the default settings for each game on the GPT-3.5 (0125) model. Utilizing this model as a case study, we illustrate our methodology for benchmarking an LLM with $\gamma$-Bench. The prompt and its design method can be found in §C in the appendix. Each game involves ten agents based on GPT-3.5, with the temperature parameter set to one. For simultaneous games, there will be twenty rounds. We run each game five times to enhance the reliability of our findings and mitigate the impact of variance. For clarity and conciseness, this section presents one of the five runs while §4.1 details quantitative results. Our findings of GPT-3.5’s behaviors on $\gamma$-Bench include:
Key Findings: The model’s decisions are mainly influenced by the outcomes of the preceding round rather than deriving from the reasoning of the optimal strategy. Although initially demonstrating suboptimal performance, the model can learn from historical data and enhance its performance over time. The model demonstrates the ability to engage in spontaneous cooperation, leading to increased social welfare beyond mere self-interest, without the necessity for explicit communication. However, this phenomenon also results in low performance in Betraying Games. The model shows limitations in sequential games characterized by more complicated rules. The aggregate score of the model on $\gamma$-Bench is $44.9$.
3.1 Cooperative Games
(1) Guess 2/3 of the Average
The vanilla setting for this game is $MIN=0$, $MAX=100$, and $R=\frac{2}{3}$. We show the choices made by all agents as well as the average and the winning numbers in Fig. 2(1). Key observations are: (1) In the first round, agents consistently select $50$ (or close to $50$), corresponding to the mean of a uniform distribution ranging from 00 to $100$. This behavior suggests that the model fails to recognize that the winning number is $\frac{2}{3}$ of the average. (2) As rounds progress, the average number selected decreases noticeably, demonstrating that agents are capable of adapting based on historical outcomes. Since the optimal strategy is to choose the $MIN$, the score in this game is given by $S_{1}=\frac{1}{NK}\sum_{ij}(C_{ij}-MIN)$, where $C_{ij}$ is the chosen number of player $i$ in round $j$. The model scores [^2] For clarity, we normalize raw scores to the range of $[0,100]$, with higher values indicating a better performance. The method used for rescaling is detailed in §E of the appendix. $65.4$ on this game.
(2) El Farol Bar
The vanilla setting for this game is $MIN=0$, $MAX=10$, $HOME=5$, and $R=60\%$. To explore the influence of incomplete information, we introduce two settings: Explicit indicates that everyone can see the results at the end of each round, while Implicit indicates that those staying at home cannot know what happened in the bar after the round ends. Fig. 2(2) illustrates the probability of agents deciding to go to the bar and the total number of players in the bar. We find that: (1) In the first round, there is an inclination among agents to visit the bar. Observations of overcrowding lead to a preference for staying home, resulting in fluctuations shown in both Fig. 2(2-1) and Fig. 2(2-2). In the Implicit setting, due to the lack of direct observations of the bar’s occupancy, agents require additional rounds (Rounds $2$ to $6$) to discern the availability of space in the bar. (2) The probability of agents going to the bar gradually stabilizes, with the average probability in the Implicit setting being lower than in the Explicit setting. Since the optimal strategy is to choose the go with a probability of $R$, the raw score [^3] For simplicity, we evaluate only the Implicit setting. in this game is given by $S_{2}=\frac{1}{K}\sum_{j}\lvert\frac{1}{N}\sum_{i}D_{ij}-R\rvert$, where $D_{ij}=1$ when player $i$ chose to go in round $j$ and $D_{ij}=0$ when player $i$ chose to stay. The model scores $73.3$ on this game.
Figure 2: Performance of GPT-3.5 (0125) in Cooperative and Betraying games.
(3) Divide the Dollar
The vanilla setting for this game is $G=100$. We plot the proposals by all agents and the sum of their proposals in Fig. 2(3). Our analysis reveals the following insights: (1) In the first round, agents’ decisions align with the NE predictions of the game. However, after gaining golds, agents exhibit increased greed, proposing allocations that exceed the NE-prescribed amounts. Upon receiving nothing, they tend to propose a “safer” amount. The trend continues and causes fluctuations across subsequent rounds. (2) Despite these fluctuations, the average of proposed golds converges to approximately $100$. Since the optimal strategy is to propose $G/N$, the raw score in this game is given by $S_{3}=\frac{1}{K}\sum_{j}\lvert\sum_{i}B_{ij}-G\rvert$, where $B_{ij}=1$ is the proposed amount number of player $i$ in round $j$. The model scores $68.1$ on this game.
3.2 Betraying Games
(4) Public Goods Game
The vanilla setting for this game is $R=2$. Each player has $T=20$ to contribute in each round. Fig. 2(4) shows the contributed tokens by each agent and their corresponding gains per round. The observations reveal the following: (1) Despite an investment return of $-80\%$, agents display a pattern of alternating between free-riding and contributing all their tokens. (2) As the rounds progress, there is an evident increase in the number of tokens contributed to the public pot, leading to an overall enhancement in social welfare gains. These findings suggest that the LLM exhibits cooperative behavior, prioritizing collective benefits over individual self-interest. Since we expect the model to infer the optimal strategy, i.e., contributing zero tokens, the raw score in this game is given by $S_{4}=\frac{1}{NK}\sum_{ij}C_{ij}$, where $C_{ij}=1$ is the proposed contribution amount of player $i$ in round $j$. The model scores $41.3$ on this game.
(5) Diner’s Dilemma
The vanilla setting for this game is $P_{h}=20$, $P_{l}=10$, $U_{h}=20$, $U_{l}=15$. We show the probability of agents choosing the costly dish, their resulting utilities, and the average bill in Fig. 2(5). Analysis of the figure reveals the following insights: (1) Contrary to the NE predictions for this game, agents predominantly prefer the cheap dish, which maximizes total social welfare. (2) Remarkably, a deviation from cooperative behavior is observed wherein one agent consistently chooses to betray others, thereby securing a higher utility. This pattern of betrayal by this agent persists across subsequent rounds. Since we expect the model to infer the the optimal strategy, i.e., choosing the costly dish, the raw score in this game is given by $S_{5}=\frac{1}{NK}\sum_{ij}D_{ij}$, where $D_{ij}=1$ when player $i$ chose the cheap dish in round $j$ and $D_{ij}=0$ when player $i$ chose the costly dish. The model scores $4.0$ on this game.
(6) Sealed-Bid Auction
For the vanilla setting in this game, we randomly assign valuations to each agent in each round, ranging from 00 to $200$. We fix the seed for random number generation to ensure fair comparisons across various settings and models. We evaluate LLMs’ performance under both First-Price and Second-Price settings. Fig. 2(6) depicts the subtraction between valuations and bids and bid amounts of each agent. Our key findings include: (1) As introduced in §2.2, we note that agents generally submit bids that are lower than their valuations in the First-Price auction, a tendency indicated by the positive discrepancies between valuations and bids depicted in Fig. 2(6-1). (2) Though the NE suggests that everyone bids the amount of their valuation in the Second-Price setting, we find a propensity for bidding below valuation levels, as demonstrated in Fig. 2(6-2). Since the optimal strategy is to bid the prices lower than their true valuations [^4] We evaluate only the First-Price setting according to the definition of Betraying Games., the raw score in this game is given by $S_{6}=\frac{1}{NK}\sum_{ij}(v_{ij}-b_{ij})$, where $v_{ij}$ and $b_{ij}$ are player $i$’s valuation and bid in round $j$, respectively. The model scores $6.5$ on this game.
3.3 Sequential Games
(7) Battle Royale
Figure 3: GPT-3.5 (0125)’s performance in “Battle Royale.” (a): Agents’ actions and outcomes of each round. For example, in round $11$, player $6$ shot at player $7$ but missed.
Pirate Rank 1 2 3 4 5 6 7 8 9 10 $S_{8P}$ $S_{8V}$
Round 1 100✓ 0✗ 0✗ 0✗ 0✗ 0✗ 0✗ 0✗ 0✗ 0✗ $8$ $1.00$
Round 2 - 99✓ 0✗ 1✓ 0✓ 0✗ 0✗ 0✗ 0✗ 0✓ $6$ $0.75$
Round 3 - - 50✓ 1✓ 1✓ 1✓ 1✓ 1✓ 1✓ 44✓ $94$ $0.57$
Table 3: Performance of GPT-3.5 (0125) in the “Pirate Game.” Each row shows the proposed gold distribution in the specific round and whether each pirate accepts (“✓”) or rejects (“✗”) the proposal. $S_{8P}$ shows the score of the proposer while $S_{8V}$ shows the score of all voters.
(8) Pirate Game
4 Further Experiments
This section explores deeper into several following Research Questions (RQs). RQ1 Robustness: Is there a significant variance in multiple runs? Is the performance sensitive to different temperatures and prompt templates? RQ2 Reasoning Strategies: Are strategies to enhance reasoning skills applicable to game scenarios? This includes implementing Chain-of-Thought (CoT) Kojima et al. (2022); Wei et al. (2022) reasoning and assigning unique personas to LLMs. RQ3 Generalizability: How does LLM performance vary with different game settings? Do LLMs remember answers learned during the training phase? RQ4 Leaderboard: How do various LLMs perform on $\gamma$-Bench? Unless otherwise specified, we apply the vanilla settings described in §3.
4.1 RQ1: Robustness
Multiple Runs
Firstly, we run all games five times under the same settings. Fig. 4 illustrates the average performance across tests, while Table 5 lists the corresponding scores. The analysis reveals that, except for the two sequential games and the “Public Goods Game,” the model demonstrates a consistent performance, as evidenced by the low variance in scores for each game.
Temperatures
As discussed in our literature review in §B, prior research incorporates varying temperature parameters from 00 to $1$ yet omits to explore their impacts. This study conducts experiments across games employing a range of temperatures $\{0.0,0.2,0.4,0.6,0.8,1.0\}$ under vanilla settings. The results, both visual and quantitative, are documented in Fig. 5 and Table 6, respectively. The small overall variance of $3.4$ indicates that, for the majority of games, temperature adjustments yield negligible effects. A notable exception is observed in “Guess 2/3 of the Average,” where increased temperatures correlate with enhanced scores ($48.0$ to $65.4$), contrasting starkly with the near-random performance at zero temperature.
Prompt Templates
We also investigate the impact of prompt phrasing on model performance. Leveraging GPT-4, we rewrite our default prompt templates described in §C, generating four additional versions. We perform a manual checking process on the generated versions to ensure GPT-4’s adherence to game rules without altering critical data. The prompt templates can be found in §D. We plot the results of using these templates in Fig. 6 and record the quantitative scores in Table 7. Notably, we find that prompt wording can significantly affect performance, as shown by the high variances in the “Public Goods Game” ($11.5$), “Diner’s Dilemma” ($23.7$), and “Pirate Game” ($14.6$).
Answer to RQ1: GPT-3.5 exhibits consistency in multiple runs and shows robustness against different temperature settings. However, inappropriate prompt designs can significantly hurt its performance.
4.2 RQ2: Reasoning Strategies
This RQ focuses on improving the model’s performance through prompt instructions. We investigate two strategies: Chain-of-Thought (CoT) prompting Kojima et al. (2022) and persona assignment Kong et al. (2023). We show the visualized and quantitative results in Fig. 7 and Table 8.
CoT
According to Kojima et al. (2022), introducing a preliminary phrase, “Let’s think step by step,” encourages the model to sequentially analyze and explain its reasoning before presenting its conclusion. This approach has proven beneficial in specific scenarios, such as games (1), (3), (4), and (5), improving the overall score from $44.9$ to $57.4$, by $12.5$. In the “(3) Divide the Dollar” game, incorporating CoT reduces the model’s propensity to suggest disproportionately large allocations, increasing the score by $15.3$. Similarly, in the “(4) Public Goods Game” and “(5) Diner’s Dilemma,” CoT prompts the model to recognize being a free-rider as the optimal strategy, increasing the scores by $14.8$ and $51.5$, respectively.
Persona
Kong et al. (2023) demonstrated that assigning specific roles to models enhances performance across various downstream tasks. Inspired by this discovery, our study initiates with a prompt that specifies the model’s role, such as “You are [ROLE],” where the role could be a cooperative and collaborative assistant, a selfish and greedy assistant, or a mathematician. Our findings reveal that assigning the “cooperative” role enhances model performance in games (1), (2), and (3), notably outperforming the CoT method in the “(3) El Farol Bar” game. Conversely, the “selfish” role markedly diminishes performance almost all the games, with the only exception of the “(7) Battle Royale” game. The “mathematician” role improves the model’s overall score by $0.8$, which is small and does not surpass the CoT method’s effectiveness.
Answer to RQ2: It is possible to improve GPT-3.5 through simple prompt instructions. Among the methods we explore, the CoT prompting performs the best, achieving a performance close to GPT-4 ($57.4$ vs. $60.5$).
4.3 RQ3: Generalizability
Considering the extensive exploration of games in domains such as mathematics, economics, and computer science, it is probable that the vanilla settings of these games are included within the training datasets of LLMs. To ascertain the presence of data contamination in our chosen games, we subjected them to various settings. The specifics of the parameters selected for each game are detailed in Table 9, and the experimental outcomes are visually represented in Fig. 8. Our findings indicate variability in model generalizability across different games. Specifically, in games (1), (5), (6), (7), and (8), the model demonstrated consistent performance under diverse settings. Conversely, the model exhibited low generalizability in games (2), (3), and (4). An analysis of the game “(2) El Farol Bar” reveals a consistent decision-making pattern by the model, opting to participate with approximately a $50\%$ probability regardless of varying bar capacities ($R$). Similarly, in the “(4) Public Goods Game,” the model consistently contributes similar amounts, even when the return rate was nil, indicating a lack of understanding of the game rules. In the game of “(3) Divide the Dollar,” the model’s performance improved with an increase in total golds ($G$), suggesting that higher allocations of golds satisfy the demands of all players.
Answer to RQ3: GPT-3.5 demonstrates variable performance across different game settings, exhibiting notably lower efficacy in “(2) El Farol Bar” and “(4) Public Goods Game.” It is noteworthy that, $\gamma$-Bench provides a test bed to evaluate the ability of LLMs in complex reasoning scenarios. As model’s ability improves (e.g., achieving more than $90$ on $\gamma$-Bench), we can increase the difficulty by varying game settings.
4.4 RQ4: Leaderboard
This RQ investigates the variance in decision-making capabilities among different LLMs, using $\gamma$-Bench. We first focus on open-source models, including OpenAI’s GPT-3.5 (0613, 1106, and 0125), GPT-4 (0125), and Google’s Gemini Pro (1.0, 1.5). The results are organized in Table 1, with model performance visualized in Fig. 9 in the appendix. Our findings reveal that Gemini-Pro (1.5) markedly surpasses other models, particularly in games (1), (4), and (5). GPT-4 follows closely behind Gemini-Pro, achieving $60.5$. The lowered performance ($23.0$) in the “(2) El Farol Bar” game stems from its conservative strategy favoring staying home. Its underperformance ($0.9$) in the “(5) Diner’s Dilemma” game is attributed to a preference for individual gain over collective benefit. Furthermore, an evaluation of three GPT-3.5 updates shows similar performance.
Next, we focus on open-source models, whose performance is detailed in Table 2 and visualized in Fig. 10. The top-performing open-source model, LLaMA-3.1-70B, closely follows Gemini-1.5-Pro with a score of $60.9$, surpassing GPT-4. Most open-source models, including Qwen-2, Mixtral-8x22B, and LLaMA-3.1-8B, outperform GPT-3.5 and Gemini-1.0-Pro. However, Mixtral-8x7B exhibits the lowest performance, likely due to its smaller size and weaker reasoning capabilities. Interestingly, LLaMA-3.1-405B underperforms compared to its smaller counterpart, the 70B version, which we attribute to its overly conservative strategy in the “(2) El Farol Bar” game, a challenge similar to the one faced by GPT-4.
Answer to RQ4: Currently, Gemini-1.5-Pro outperforms all other models evaluated in this study. LLaMA-3.1-70B performs closely, being in the second place.
5 Related Work
5.1 Specific Games
5.2 Game Benchmarks
6 Conclusion
This paper presents $\gamma$-Bench, a benchmark designed to assess LLMs’ Gaming Ability in Multi-Agent environments. $\gamma$-Bench incorporates eight classic game theory scenarios, emphasizing multi-player interactions across multiple rounds and actions. Our findings reveal that GPT-3.5 (0125) demonstrates a limited decision-making ability on $\gamma$-Bench, yet it can improve itself by learning from the historical results. Leveraging the carefully designed scoring scheme, we observe that GPT-3.5 (0125) exhibits commendable robustness across various temperatures and prompts. It is noteworthy that strategies such as CoT prove effective in this context. Nevertheless, its capability to generalize across various game settings remains restricted. Finally, Gemini-1.5-Pro outperforms all tested models, achieving the highest ranking on the $\gamma$-Bench leaderboard, with the open-source LLaMA-3.1-70B following closely behind.
Appendix A More Information on Game Theory
A.1 Formulation